Two large vertical and parallel metal plates having a separation of 1 cm are connected to a DC voltage source of potential difference X. A proton is released at rest midway between 

the two plates. It is found to move at 45º to the vertical JUST after release. Then X is 

nearly 

(A) 1* 106^ -5V

(B) 1* 10^- 7V

(C) 1* 10^- 9V

(D) 1 *10^- 10V 

Grade:12

1 Answers

Mukesh Sharma
10 years ago
As we know that F=qE also F=mg due to mass and gravitation
there fore qE=mg
q(v/d)=mg
V=mgd/q
now putting the values v=1.67*10-27*10*10-2/1.6*10-19
we get V=10-9 V hence the option (c) is correct

Thanks & regards

Mukesh Sharma

Ask iiTians Faculty

M.sc, Jiwaji university , Gwalior